How do i solve 2x +7y =8 and x- 9y =4 ?
Answer by jim_thompson5910(35256)

Hint:

x- 9y =4

x = 9y+4

-----------------------

2x +7y =8

2x +7y =8

2(x) + 7y = 8

2(9y+4)+7y = 8

2(9y+4) + 7y = 8